OpenStudy (anonymous):

Given the equation y-3 =1\2(x+6) in point-slope form, identify the equation of the same line in standard form A. x - 2 = -12 B. y= 1\2x C. y = 1\2x + 6 D. y = 1\2x + 9

OpenStudy (texaschic101):

y - 3 = 1/2(x + 6) first thing you need to do is distribute the 1/2 through the parenthesis....can you do that ? wait....lol....this is simpler then I thought. ( that is if the first answer choice has a y after the -2) standard form is Ax + By = C.....there is only 1 answer choice in this form. Any thoughts ?

OpenStudy (anonymous):

my choice is A. I am bad at math so, coming hear is the best place for me to find the answers i need. :)

OpenStudy (texaschic101):

It would be A...but only if there was a y after the -2

OpenStudy (anonymous):

i did not read the first part of your reply :/. i am so sorry :/

OpenStudy (texaschic101):

unless the question is asking for slope intercept form ??

OpenStudy (anonymous):

i am going with C

OpenStudy (texaschic101):

no...wait....let me show you y - 3 = 1/2(x + 6) y - 3 = 1/2x + 3 y = 1/2x + 3 + 3 y = 1/2x + 6 <=== slope intercept form y = 1/2x + 6 -- subtract 1/2x -1/2x + y = 6 -- multiply by -2 x - 2y = -12 <== standard form

OpenStudy (anonymous):

so...... its A?

OpenStudy (texaschic101):

if A, the way you posted in your answer choices has no y in it.....if it is supposed to be : x - 2y = -12, then that is the correct answer if you need it in standard form

OpenStudy (anonymous):

i forgot the y. *slaps hand on forhead*

OpenStudy (texaschic101):

lol....its A then

OpenStudy (anonymous):

thank you so much!!!! your the best!!!!

OpenStudy (texaschic101):

lol...not sure about that...but thanks

OpenStudy (anonymous):

can you help me with another question?

OpenStudy (texaschic101):

I can try...can't make no promises

OpenStudy (anonymous):

choose the point-slope form of the equation below that represents the line that passes through the points (-3,2) and (2,1) A. y+3 = -5(x,2) B. y-2 = -5(x+3) C. y+3 =-1/5(x-2) D. y-2 = -1/5(x+3)

OpenStudy (texaschic101):

first we find the slope by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) now we have 2 sets of points... set 1 = (-3,2)....so we will call these x1 and y1 --- (x1 = -3 and y1 = 2) set 2 = (2,1)...so we will call these x2 and y2 --- (x2 = 2 and y2 = 1) now we sub into our slope formula slope(m) = (y2 - y1) / (x2 - x1) slope(m) = (1 - 2) / (2 - (-3) slope(m) = -1/(2 + 3) slope(m) = -1/5 now we need it in point-slope form : y - y1 = m(x - x1) slope(m) = -1/5...so that automatically eliminates A and B lets use either of our points (-3,2)....x1 = -3 and y1 = 2 so now we sub into our point-slope formula y - y1 = m(x - x1) y - 2 = -1/5(x - (-3) y - 2 = -1/5(x + 3) <=== this is your point-slope form questions ??

OpenStudy (anonymous):

so its D? i suck at math, but i have you to help me :)

OpenStudy (texaschic101):

yes...its D

OpenStudy (anonymous):

thank you, thank you, thank you!!!!!!!!!!!!

OpenStudy (texaschic101):

YW :)

OpenStudy (anonymous):

this next question is the same as the other one, but different numbers

OpenStudy (texaschic101):

ok...you want to try to do it and then let me check your work ?

OpenStudy (anonymous):

sure....................................... :)

OpenStudy (texaschic101):

tag me when your ready for me to check your work...or if you are having a problem

OpenStudy (anonymous):

i got B

OpenStudy (texaschic101):

kinda need to see the problem